OpenStudy (unofficialllyy):

A parachutist's rate during a free fall reaches 198 feet per second. What is this rate in meters per second? At this rate, how many meters will the parachutist fall during 20 seconds of free fall? In your computations, assume that 1 meter is equal to 3.3 feet. Do not round your answers.

OpenStudy (unofficialllyy):

I have to find the rate and the distance fallen in 20 seconds

OpenStudy (jhannybean):

\[\sf \frac{198~ft}{1~sec}~\times \frac{1~m}{3.3~ft} = ~\frac{198}{3.3}~\frac{m}{s} \]Now create a ratio. \[\sf \frac{198~m}{3.3~sec} = \frac{x~m}{20~sec}\] Cross multiple and solve for the meters fallen in 20 seconds.

OpenStudy (unofficialllyy):

Thank you I got it :)

OpenStudy (jhannybean):

nice :)